# StreamingThinker: Large Language Models Can Think While Reading

### v2

@@ Appendix D Dataset Details
We construct math reasoning data from GSM-Symbolic (mirzadehgsm), MetaMathQA (yumetamath), and TuLu-personas-math-grade (lambert2024tulu).
@@
For logical reasoning, we evaluate on ProofWriter (tafjord2020proofwriter) and LogicNLI (tian2021diagnosing). ProofWriter is a benchmark designed for evaluating multi-step logical reasoning and theorem proving capabilities in natural language. It is derived from the EntailmentBank corpus, where each example consists of a hypothesis and a set of supporting facts expressed in natural language. The task requires the model to infer whether the hypothesis is entailed, contradicted, or neutral given the supporting facts, while optionally generating an explicit reasoning chain that connects the premises to the conclusion. LogicNLI is a natural language inference (NLI) benchmark designed to evaluate the logical reasoning ability of language models beyond surface-level semantics. Each example in LogicNLI consists of a premise and a hypothesis pair, annotated with one of three logical relations: entailment, contradiction, or neutral. Unlike conventional NLI datasets that focus primarily on lexical or syntactic cues, LogicNLI emphasizes reasoning over formal logic structures such as conjunction, disjunction, negation, quantifiers, and implication.
@@
For context-based QA reasoning task, we evaluate StreamingThinker on PubMedQA (jin2019pubmedqa) and HotpotQA (yang2018hotpotqa) datasets.
@@ E.3 Relation between Parallel KV Caches and Prefill-Decode Separation
While our parallel KV caches mechanism may formally resemble techniques for Prefill-Decode separation (zhong2024distserve; qin2024mooncake), the foundational motivations and operational goals behind the two approaches are fundamentally distinct.

### v3

@@ Appendix D Dataset Details
We construct math reasoning data from GSM-Symbolic (Mirzadeh et al., 2025), MetaMathQA (Yu et al., 2024), and TuLu-personas-math-grade (Lambert et al., 2024).
@@
For logical reasoning, we evaluate on ProofWriter (Tafjord et al., 2020) and LogicNLI (Tian et al., 2021). ProofWriter is a benchmark designed for evaluating multi-step logical reasoning and theorem proving capabilities in natural language. It is derived from the EntailmentBank corpus, where each example consists of a hypothesis and a set of supporting facts expressed in natural language. The task requires the model to infer whether the hypothesis is entailed, contradicted, or neutral given the supporting facts, while optionally generating an explicit reasoning chain that connects the premises to the conclusion. LogicNLI is a natural language inference (NLI) benchmark designed to evaluate the logical reasoning ability of language models beyond surface-level semantics. Each example in LogicNLI consists of a premise and a hypothesis pair, annotated with one of three logical relations: entailment, contradiction, or neutral. Unlike conventional NLI datasets that focus primarily on lexical or syntactic cues, LogicNLI emphasizes reasoning over formal logic structures such as conjunction, disjunction, negation, quantifiers, and implication.
@@
For context-based QA reasoning task, we evaluate StreamingThinker on PubMedQA (Jin et al., 2019) and HotpotQA (Yang et al., 2018) datasets.
@@ E.3 Relation between Parallel KV Caches and Prefill-Decode Separation
While our parallel KV caches mechanism may formally resemble techniques for Prefill-Decode separation (Zhong et al., 2024; Qin et al., 2024), the foundational motivations and operational goals behind the two approaches are fundamentally distinct.